# Multimodal Industrial Anomaly Detection by Crossmodal Reverse Distillation

### v2

@@ Abstract
Knowledge distillation (KD) has been widely studied in unsupervised image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored. Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges. Anomalies in one modality may not be effectively captured in the fused teacher features, leading to detection failures. Besides, these methods do not fully leverage the rich intra- and inter-modality information. In this paper, we propose Crossmodal Reverse Distillation (CRD) based on Multi-branch design to realize Multimodal Industrial AD. By assigning independent branches to each modality, our method enables finer detection of anomalies within each modality. Furthermore, we enhance the interaction between modalities during the distillation process by designing Crossmodal Filter and Amplifier. With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected. Experimental verifications on multimodal AD datasets demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization. Code will be available.
1 Introduction
@@
Figure 1: Qualitative results on MVTec 3D-AD [7]. From top to bottom: 2D images, 3D depth maps, the ground truth masks, the output anomaly maps of multimodal AD methods based on feature learning showing AST [26] with more false negatives, Crossmodal Feature Mapping [12] (abbreviated as CFM) with more false positives, and our method CRD with superior segmentation.
Figure 2: Comparison of our proposed CRD with other multimodal AD methods based on feature learning.
@@
Early unsupervised multimodal AD methods often relies on memory banks, such as BTF [20] and M3DM [31]. These methods construct large memory banks by processing and storing multimodal features of normal samples. During inference, extracted features are compared with normal features in the memory bank to determine if these samples are anomalous. However, such methods inevitably require substantial storage, which becomes a bottleneck, especially when dealing with large-scale datasets.
@@
Recent studies introduce the idea of feature learning to multimodal AD to avoid large-scale storage [26, 15, 12]. The feature learning methods typically leverage pretrained networks which is frozen during training to generate rich feature representations and transfer them to trainable networks. Since only normal samples are used during training, the trainable networks only learn the normality representation ability from the frozen networks. In other words, the abnormality representation ability is not transferred. During inference, the distances between features from the frozen and trainable networks are used to detect anomalies.
@@
Currently, feature learning-based methods are divided into two categories as in Fig. 2: (1) Methods like AST [26] and MMRD [15] (Fig. 2 (a)) follow Knowledge Distillation, where student networks learn the multimodal feature representation ability from teacher networks on normal samples. However, these methods may smooth out abnormal features during fusion, reducing the teacher’s sensitivity to anomalies and leading to false negatives, particularly when one modality is normal and the other is anomalous. (2) Crossmodal Feature Mapping [12] (Fig. 2 (b)) introduces a new paradigm, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships. But inevitably, modal discrepancy between 2D and 3D features results in underfitting and causes false positives in the output anomaly maps.
@@
To address the anomaly smoothing issue in previous KD-based methods, our intuition is to expand the single-branch distillation into a multi-branch one. Each modality has its own distillation target, ensuring the sensitivity to anomalies within each modality. Additionally, to tackle the underfitting problem of Crossmodal Feature Mapping, we propose not to directly use the feature similarity from crossmodal mapping for anomaly map generation. Instead, we integrate the idea of crossmodal mapping into the multi-branch distillation, where features from another modality’s teacher network, along with the inter-modal relationship, help generate features for the student network of a given modality.
@@
In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD) [13], named Crossmodal Reverse Distillation (CRD). CRD naturally combines the strengths of both feature learning paradigms. First, we design Multi-branch Distillation that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps. Next, we introduce two crossmodal assistants within each modality branch. One is Crossmodal Filter, which uses information from other modalities to help the student decoder generate normal features. And the other is Crossmodal Amplifier, which allows a modality branch to detect anomalies from other modalities. Finally, anomaly scores of test samples are calculated by integrating anomaly maps of all branches. Experimental results on representative multimodal AD datasets demonstrate the effectiveness of CRD in multimodal anomaly detection and localization.
2 Related Work
@@
Unsupervised Anomaly Detection. Unsupervised AD has gained increasing attention in recent years. Initially, many unsupervised AD methods relied on generative models [5, 1, 38, 32, 19, 18]. These models are trained on normal samples to learn the ability to reconstruct normal data. During inference, reconstruction errors are used to classify input images as normal or anomalous. Other methods introduce memory banks [25, 2, 22], comparing test samples with stored normal features to detect anomalies. Recently, significant progress has been made in the research of synthetic anomaly images [21, 34, 23], which simulate real-world scenarios to assist in unsupervised AD tasks. Additionally, knowledge distillation methods [27, 30, 26, 3, 13, 28, 16, 14, 37, 17] based on the teacher-student framework have also been applied to unsupervised AD. These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies. Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
@@
Multimodal Anomaly Detection. With the release of multiple 3D industrial anomaly detection datasets [7, 9, 24], unsupervised multimodal AD has gradually become a topic of research. Some methods [11, 29], such as BTF [20] and M3DM [31], follow PatchCore [25] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method. Other methods [10, 35, 36] rely on reconstruction networks to detect anomalies based on the reconstruction results of multimodal data. Knowledge distillation paradigms have also been explored in some methods, such as 3D-ST [4], AST [26], and MMRD [15], with the goal of having the student network mimic the multimodal features output by the teacher network. Crossmodal Feature Mapping [12] proposes a new multimodal solution, using two neural networks to perform crossmodal mapping and locate anomalies based on the mapping results.
3 Preliminaries
@@
4 Crossmodal Reverse Distillation
Figure 3: Pipeline of our proposed method. The feature evolution visualization highlights that anomalies in each modality are effectively captured. In resulting anomaly maps, darker gray regions indicate higher anomaly scores.
Figure 4: Information flow of the branch of Modality A, showing how Crossmodal Filter and Crossmodal Amplifier incorporate information from Modality B. Crossmodal Filter aids the student decoder in generating normal features by utilizing features from Modality B, while Crossmodal Amplifier strengthens the perception of anomalies in Modality A by amplifying anomaly signals from Modality B.
In the multimodal AD task, each input consists of multiple modalities of data. In this section, we primarily focus on 2D (e.g., RGB images) and 3D (e.g., depth maps) modalities as examples for studying multimodal anomaly detection. Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{2D},\mathcal{I}^{3D}\}$, where $\mathcal{I}^{2D}$ and $\mathcal{I}^{3D}$ are the 2D and 3D modalities respectively. The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$. The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which includes both normal and anomalous samples.
4.1 Multi-branch Distillation
In multimodal industrial AD, several methods have incorporated knowledge distillation, which are broadly classified into two categories: (1) Some methods add or concatenate multimodal data together at the input stage, transforming the multi-modal into a single-modal problem [26]. (2) Others employ multiple teacher networks to independently extract features from different modalities, and then merges the features as the learning target for the distillation of student network [15]. The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern. During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features from the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
@@ 4.1 Multi-branch Distillation
To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig. 3, which differs from previous KD-based multimodal AD methods with a multi-branch design. (In this section, we focus on the example case of dual-branch including 2D and 3D branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality. During training, the pre-trained teacher network independently distills knowledge for each modality. At inference, an anomaly map is generated for each modality, and these maps are fused after normalization. By deferring the fusion of modalities to the inference stage, our method MBD allows for more precise integration of multimodal information, effectively preventing the smoothing of anomalies during feature fusion and reducing the risk of false negatives.
@@
For the design of each branch, we draw inspiration from RD. Both 2D and 3D branches consist of four components: the frozen pre-trained teacher encoder, the trainable student decoder, and two crossmodal assistants including Crossmodal Filter and Crossmodal Amplifier (Sec. 4.2). The teacher encoders in both branches share parameters and use a WideResNet-50 pre-trained on ImageNet. The student decoders follow RD, using a structure symmetric to that of the teacher networks. Crossmodal Filters replace the OCBE module of RD, while Crossmodal Amplifiers are placed after the student decoders to optimize the output features, enabling crossmodal mapping and information fusion between modalities. The anomaly map for each branch is obtained by calculating the cosine distance between the output features of Crossmodal Amplifier and the teacher encoder.
4.2 Crossmodal Assistants
@@
To fully leverage the complementary information between modalities, we introduce crossmodal interactions within Multi-branch Distillation. To be specific, we design two crossmodal assistants serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies. (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality. With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final anomaly map.
@@
Crossmodal Filter. The training objective of RD is to ensure that the student decoder generates anomaly-free features. It is achieved through the OCBE bottleneck module in vanilla RD. By further compressing the encoder features, the OCBE module effectively prevents anomaly disturbances from propagating into the student decoder, ensuring that the student decoder generates anomaly-free features even when anomaly samples are input during inference. However, in multimodal AD, relying only on the OCBE module to get the student input results in underutilization of the rich multimodal information. In particular, when one modality is anomalous and the other is normal, we believe that using the information from the normal modality aids in reconstructing the anomaly-free features of the anomalous modality. Therefore, we propose Crossmodal Filter to optimize the OCBE process by incorporating the normal information from the other modality.
@@
It is worth noting that to ensure the features of modality B used for assistance are normal, we mimic the concept of OCBE and design Bottleneck Projection as a two-step process of compression and restoration. First, the three encoder features ${F^{B}_{E}}^{i},i=\{1,2,3\}$ of modality B are downsampled to $8\times 8$ by convolution. Then, these features are upsampled back to their original size using deconvolution to obtain ${F^{B}_{BP}}^{i},i=\{1,2,3\}$. During training, we minimize the distance between $F^{B}_{BP}$ and the teacher encoder features of modality A $F^{A}_{E}$, mapping modality B’s features into modality A’s feature space. During inference, the feature compact process of BP helps reduce the interference of anomalies in modality B, ensuring that normal information from modality B is effectively transferred to the modality A’s branch.
Crossmodal Amplifier. When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Sec. 4.1. We address this issue with Multi-branch Distillation. However, fusion of the anomaly maps still cause the anomaly values to become less noticeable after summation. Our idea is that multimodal information interaction helps resolve this issue. Specifically, when modality A is normal and modality B is anomalous, if the anomaly information from modality B is incorporated into modality A’s branch, the anomaly map generated by modality A is also able to show the anomalies in modality B. This allows the anomaly in modality B to be emphasized in modality A’s branch, thus preventing the anomaly smoothing issue during anomaly map fusion.
As illustrated in Fig. 4, we design Crossmodal Amplifier. First, Inverted Bottleneck Projection (IBP) is designed to map modality B’s features to modality A’s feature space, yielding ${F^{B}_{IBP}}^{i},i=\{1,2,3\}$. Then, the mapped features are then dynamically fused with the features ${F^{A}_{D}}^{i},i=\{1,2,3\}$ which are reconstructed by the student decoder in modality A’s branch, to obtain ${F^{A}_{CA}}^{i},i=\{1,2,3\}$ by

### v3

# Tuned Reverse Distillation: Enhancing Multimodal Industrial Anomaly Detection with Crossmodal Tuners

@@ Abstract
Knowledge distillation (KD) has been widely studied in unsupervised image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored. Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges. Anomalies that only exist in one modality may not be effectively captured in the fused teacher features, leading to detection failures. Besides, these methods do not fully leverage the rich intra- and inter-modality information that are critical for effective anomaly detection. In this paper, we propose Tuned Reverse Distillation (TRD) based on Multi-branch design to realize Multimodal Industrial AD. By assigning independent branches to each modality, our method enables finer detection of anomalies within each modality. Furthermore, we enhance the interaction between modalities during the distillation process by designing two Crossmodal Tuners including Crossmodal Filter and Amplifier. With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected. Experimental verifications on multimodal AD datasets demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization. Code is available at https://github.com/hito2448/TRD.
Index Terms:
Multimodal anomaly detection, knowledge distillation, unsupervised anomaly detection, anomaly localization.
I Introduction
@@
Figure 1: Qualitative comparisons on MVTec 3D-AD [4]. From top to bottom: 2D images, 3D depth maps, the ground truth masks, the output anomaly maps of multimodal AD methods based on feature learning showing AST [5] with more false negatives, Crossmodal Feature Mapping [6] (abbreviated as CFM) with more false positives, and our method TRD with superior segmentation.
@@
Early unsupervised multimodal AD methods often rely on memory banks, such as BTF [7] and M3DM [8]. These methods construct large memory banks by processing and storing multimodal features of normal samples. During inference, extracted features are compared with normal features in the memory bank to determine if these samples are anomalous. However, such methods inevitably require substantial storage, which becomes a bottleneck, especially when dealing with large-scale datasets.
@@
Recent studies introduce the idea of feature learning to multimodal AD to avoid large-scale storage [5, 9, 6]. The feature learning-based methods typically leverage pre-trained networks which is frozen during training to generate rich feature representations and transfer them to trainable networks. Since only normal samples are used during training, the trainable networks only learn the normality representation ability from the frozen networks. In other words, the abnormality representation ability is not transferred. During inference, the distances between features from the frozen and trainable networks are used to detect anomalies.
Figure 2: Comparison of our proposed TRD with previous KD-based methods. Previous KD-based methods tend to suffer from anomaly smoothing when one modality is normal and the other contains anomalies. For example, in this figure, the anomaly on the bagel is barely visible in the 2D modality and resembles normal samples, while it is clearly distinguishable in the 3D modality. However, one of previous KD-based methods AST [5], which fuses the two modalities before distillation, fails to detect the anomaly and results in false negative, as in (a). In contrast, our TRD employs a multi-branch design and Crossmodal Tuners to preserve modality-specific anomaly cues, enabling more accurate anomaly localization as in (b).
@@
Currently, feature learning-based methods for multimodal AD are primarily divided into two categories. One is Crossmodal Feature Mapping [6], which introduces a novel paradigm for addressing multimodal AD, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships. When anomalies occur, the failure of feature mapping enables anomaly localization. But inevitably, modal discrepancy between 2D and 3D features results in underfitting and causes false positives in the output anomaly maps.
Knowledge distillation-based methods [5, 9] represent another widely accepted category of feature learning-based methods, which extend the typical KD paradigm, originally used in image anomaly detection, to multimodal settings. In such methods, student networks are trained on normal samples to learn multimodal feature representations from teacher networks. Anomalies are then identified by measuring the feature discrepancies between the student and the teacher. To obtain multimodal teacher features as distillation targets, KD-based multimodal AD methods typically adopt one of two fusion strategies. Some fuse 2D and 3D data at the input level before passing it through a pre-trained encoder. Others extract features from separately frozen 2D and 3D encoders and then merge them. However, both fusion strategies smooth out abnormal features during fusion, reducing the teacher’s sensitivity to anomalies, particularly when one modality is normal and the other is anomalous. In such cases, the fused teacher features tend to be dominated by the normal modality, reducing their sensitivity to subtle anomalies and increasing the risk of false negatives, as illustrated in Fig. 2 (a).
@@
To address the anomaly smoothing issue in previous KD-based methods, our intuition is to expand the single-branch distillation into a multi-branch one. Each modality is assigned an independent distillation objective, allowing the student network to learn from the corresponding teacher features. This multi-branch design preserves the anomaly sensitivity of each modality and avoids the anomaly smoothing problem caused by early feature fusion. With this setup, each branch is able to accurately detect anomalies specific to its modality. However, simply performing late fusion by combining anomaly maps from each branch may still weaken anomaly signals, particularly when anomalies only in one modality are subtle. To further mitigate this issue, we integrate the idea of crossmodal mapping into the multi-branch distillation, where features from another modality’s teacher network, along with the inter-modal relationship, help generate features for the student network of a given modality. This enables each branch not only to retain its own anomaly information but also to capture crossmodal cues, ensuring that anomalies present in any modality are effectively reflected in the final output.
@@
In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD) [10], named Tuned Reverse Distillation (TRD). TRD naturally combines the strengths of existing feature learning paradigms, as shown in Fig. 2 (b). First, we design Multi-branch Distillation that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps. Next, we introduce two Crossmodal Tuners within each modality branch. One is Crossmodal Filter, which uses information from other modalities to help the student decoder generate normal features. And the other is Crossmodal Amplifier, which allows a modality branch to detect anomalies from other modalities. Finally, anomaly scores of test samples are calculated by integrating anomaly maps of all branches. Experimental results on representative multimodal AD datasets demonstrate the effectiveness of TRD in multimodal anomaly detection and localization.
Our contributions are as follows:
  • We propose Tuned Reverse Distillation based on Multi-branch Distillation for multimodal industrial AD, which effectively detects anomalies in all modalities.
  • Two Crossmodal Tuners are introduced, including Crossmodal Filter that helps the student decoder generate normal features, and Crossmodal Amplifier that amplifies the perception of anomalies from other modalities.
  • Experiments show that our proposed TRD achieves state-of-the-art performance on multimodal AD datasets.
II Related Work
II-A Unsupervised Anomaly Detection
@@
Unsupervised AD has gained increasing attention in recent years. Initially, many unsupervised AD methods relied on generative models [11, 12, 13, 14, 15, 16, 17]. These models are trained on normal samples to learn the ability to reconstruct normal data. During inference, reconstruction errors are used to classify input images as normal or anomalous. Other methods introduce memory banks [18, 19, 20], comparing test samples with stored normal features to detect anomalies. Recently, significant progress has been made in the research of synthetic anomaly images [21, 22, 23], which simulate real-world scenarios to assist in unsupervised AD tasks. Additionally, knowledge distillation methods [24, 5, 10, 25, 26, 27, 28, 29] based on the teacher-student framework have also been applied to unsupervised AD. These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies. Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
II-B Multimodal Anomaly Detection
@@
With the release of multiple 3D industrial anomaly detection datasets [4, 30, 31], unsupervised multimodal AD has gradually become a topic of research. Some methods [32, 33], such as BTF [7] and M3DM [8], follow PatchCore [18] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method. Other methods [34, 35, 36] rely on reconstruction networks to detect anomalies based on the reconstruction results of multimodal data. Knowledge distillation paradigms have also been explored in some methods, such as 3D-ST [37], AST [5], and MMRD [9], with the goal of having the student network mimic the multimodal features output by the teacher network. Crossmodal Feature Mapping [6] proposes a new multimodal solution, using two neural networks to perform crossmodal mapping and locate anomalies based on the mapping results.
III Preliminaries
@@
IV Tuned Distillation
Figure 3: Information flow of the branch for Modality A during training. This branch consists of a frozen teacher encoder, a trainable student decoder, and two Crossmodal Tuners including Crossmodal Filter and Crossmodal Amplifier. Crossmodal Filter incorporates mapped features from Modality B to guide the student decoder of Modality A in generating anomaly-free features. Crossmodal Amplifier injects anomaly signals from Modality B after the decoder, enhancing the anomaly sensitivity of Modality A.
In the multimodal AD task, each input consists of multiple modalities of data. In this section, we primarily focus on 2D (e.g., RGB images) and 3D (e.g., depth maps) modalities as examples for studying multimodal anomaly detection. Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{2D},\mathcal{I}^{3D}\}$, where $\mathcal{I}^{2D}$ and $\mathcal{I}^{3D}$ are the 2D and 3D modalities respectively. The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$. The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which include both normal and anomalous samples.
IV-A Multi-branch Distillation
In multimodal industrial AD, several methods have incorporated Knowledge Distillation, which are broadly classified into two categories:
  • Some methods add or concatenate multimodal data at the input stage, transforming the multimodal into a singlemodal problem [5].
  • Others employ multiple teacher networks to independently extract features from different modalities, and then merge the features as the learning target for the distillation of the student network [9].
The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern. During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features of the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
@@ IV-A Multi-branch Distillation
To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig. 2 (b), which differs from previous KD-based multimodal AD methods with a multi-branch design. (In this section, we focus on the example case of dual-branch including 2D and 3D branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality. During training, the pre-trained teacher network independently distills knowledge for each modality. At inference, an anomaly map is generated for each modality, and these maps are fused after normalization. By deferring the fusion of modalities to the inference stage, our method MBD allows for more precise integration of multimodal information, effectively preventing the smoothing of anomalies during feature fusion and reducing the risk of false negatives.
@@
For the design of each branch, we draw inspiration from RD. The two branches corresponding to the 2D and 3D modalities are designed symmetrically. Fig. 3 depicts the detailed architecture of one branch (for modality A), which includes four components: a frozen pre-trained teacher encoder, a trainable student decoder, and two Crossmodal Tuners including Crossmodal Filter and Crossmodal Amplifier (Section IV-B). The teacher encoder in both branches share parameters and use a WideResNet-50 pre-trained on ImageNet. The student decoders in both branches follow RD, using a structure symmetric to that of the teacher networks. Crossmodal Filters replace the OCBE module of RD, while Crossmodal Amplifiers are placed after the student decoders to tune the output features, enabling crossmodal mapping and information fusion between modalities.
Figure 4: Illustration of the Crossmodal Filter within the branch of Modality A, which consists of two components including Bottleneck Projection and OCBE. First, the encoder features from the other modality (Modality B) ${F_{E}^{B}}^{i},i=\{1,2,3\}$ are processed by the Bottleneck Projection to suppress anomaly signals and generate compact, normal-like features. These features are then concatenated with the encoder features from Modality A ${F_{E}^{A}}^{i},i=\{1,2,3\}$ and passed into the OCBE module. The resulting fused representation $F_{CF}^{A}$ retains rich multimodal information while remaining anomaly-free, and is used to guide the student decoder in reconstructing anomaly-free features within the Modality A branch.
IV-B Crossmodal Tuners
@@
To fully leverage the complementary information between modalities, we introduce crossmodal interactions within Multi-branch Distillation. To be specific, we design two Crossmodal Tuners serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies. (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality. With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final anomaly map.
IV-B1 Crossmodal Filter
@@
The training objective of RD is to ensure that the student decoder generates anomaly-free features. It is achieved through the OCBE bottleneck module in vanilla RD. By further compressing the encoder features, the OCBE module effectively prevents anomaly disturbances from propagating into the student decoder, ensuring that the student decoder generates anomaly-free features even when anomaly samples are input during inference. However, in multimodal AD, relying only on the OCBE module to get the student input results in underutilization of the rich multimodal information. In particular, when one modality is anomalous and the other is normal, we believe that using the information from the normal modality aids in reconstructing the anomaly-free features of the anomalous modality. Therefore, we propose Crossmodal Filter to optimize the OCBE process by incorporating the normal information from the other modality.
@@
It is worth noting that to ensure the features of modality B used for assistance are normal, we mimic the concept of OCBE and design Bottleneck Projection as a two-step process of compression and restoration. First, the three encoder features ${F^{B}_{E}}^{i},i=\{1,2,3\}$ of modality B are downsampled to $\times 8$ 8 by convolution. Then, these features are upsampled back to their original size using deconvolution to obtain ${F^{B}_{BP}}^{i},i=\{1,2,3\}$. During training, we minimize the distance between $F^{B}_{BP}$ and the teacher encoder features of modality A $F^{A}_{E}$, mapping modality B’s features into modality A’s feature space. During inference, the feature compact process of BP helps reduce the interference of anomalies in modality B, ensuring that normal information from modality B is effectively transferred to the modality A’s branch.
Figure 5: Illustration of the Crossmodal Amplifier within the branch of Modality A. The module is primarily composed of an Inverted Bottleneck Projection. Encoder features ${F_{E}^{B}}^{i},i=\{1,2,3\}$ from the other modality (Modality B) are first processed by this module to generate anomaly-aware features. These features are designed to resemble the encoder features of Modality A ${F_{E}^{A}}^{i},i=\{1,2,3\}$ in normal regions of Modality B, while exhibiting discrepancies in anomalous regions. This contrast enables the anomaly information from Modality B to be reflected in the Modality A branch. Finally, the amplified features are added to the student decoder output of Modality A to produce the final outputs ${F_{CA}^{B}}^{i},i=\{1,2,3\}$.
IV-B2 Crossmodal Amplifier
When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Section IV-A. We address this issue with Multi-branch Distillation. However, fusion of the anomaly maps still cause the anomaly values to become less noticeable after summation. Our idea is that multimodal information interaction helps resolve this issue. Specifically, when modality A is normal and modality B is anomalous, if the anomaly information from modality B is incorporated into modality A’s branch, the anomaly map generated by modality A is also able to show the anomalies in modality B. This allows the anomaly in modality B to be emphasized in modality A’s branch, thus preventing the anomaly smoothing issue during anomaly map fusion.
As illustrated in Fig. 5, we design Crossmodal Amplifier. First, Inverted Bottleneck Projection (IBP) is proposed to map modality B’s features to modality A’s feature space, yielding ${F^{B}_{IBP}}^{i},i=\{1,2,3\}$. Then, the mapped features are dynamically fused with the features ${F^{A}_{D}}^{i},i=\{1,2,3\}$ which are reconstructed by the student decoder in modality A’s branch, to obtain ${F^{A}_{CA}}^{i},i=\{1,2,3\}$ by